OpenStudy (anonymous):

solve (y-1)dx + (x-2)dy = 0

OpenStudy (mathmate):

This is a separable DE, meaning that we can put x on one side and y on the other.

OpenStudy (anonymous):

i have solved it but don't know whether it is correct or not

OpenStudy (mathmate):

All you'd have to do is to integrate both sides. Feel free to post your answer for checking!

OpenStudy (anonymous):

( dy(x))/( dx) = (-y(x)+1)/(x-2) Divide both sides by 1-y(x): (( dy(x))/( dx))/(-y(x)+1) = 1/(x-2) Integrate wrt x integral (( dy(x))/( dx))/(-y(x)+1) dx = integral 1/(x-2) dx -log(-y(x)+1) = log(x-2)+c, where c is a constant. then y(x) = (e^(-c)-x+2)/(-x+2) answer i am getting is y(x) = (-x+c)/(-x+2)

OpenStudy (anonymous):

i am also having trouble with (x+2u+1)u'=x+2u+4 , let x+2u=y

OpenStudy (anonymous):

can u check this one also @ganeshie8 plz

OpenStudy (anonymous):

is the correct @ganeshie8 @mathmate

OpenStudy (mathmate):

I usually derive it this way: Given \((y-1)dx + (x-2)dy = 0\) transpose first term to the right, change signs, and cross multiply: \(\large \frac{dy}{y-1}=-\frac{dx}{x-2}\)

OpenStudy (anonymous):

IS THE PREVIOUS ONE CORRECT @mathmate N ALSO IN THIS QUESTION THEY HAVE MENTIONED X+2U=Y SO AM STUCK WITH THAT

OpenStudy (mathmate):

Yes, I put your solution into the original equation and it checks out (i.e. =0). Good job!

OpenStudy (anonymous):

integrating w.r.t x: we get, integral (( dy(x))/( dx))/(y(x)-1) dx = integral -1/(x-2) dx by evaluating (y(x)(x))/(y(x)-1) = -log(-x+2)+c the we get | y(x) = (-log(-x+2)+c+y(x)(x))/(-log(-x+2)+c) but what should i do with X+2U=Y consideration will the answer change with that

OpenStudy (mathmate):

(x+2u+1)u'=x+2u+4 , let x+2u=y Then u'=(y'-1)/2 (y+1)(y'-1)/2=y+4 (y+1)y'-y-1=2(y+4) (y+1)y'=3(y+3)

OpenStudy (anonymous):

thanks @mathmate

OpenStudy (anonymous):

i have one last question solve xu'=(u-x)^3+u u(1)=2 take v=u-x

OpenStudy (mathmate):

I was thinking, it might be even simpler to put y=x+2u+1

OpenStudy (anonymous):

so when i solve is it like i complete with the solution which i have posted first then apply x+2u=y then u'=(y'-1)/2 (y+1)(y'-1)/2=y+4 (y+1)y'-y-1=2(y+4) (y+1)y'=3(y+3) is that right

OpenStudy (mathmate):

It would be a similar process of substitution: xu'=(u-x)^3+u u(1)=2 take v=u-x u'=v'+1 x(v'+1)=v^3+v+x xv'=v^3+v Can you take it from here?

OpenStudy (anonymous):

can u join us @zepdrix

OpenStudy (anonymous):

solve xu'=(u-x)^3+u u(1)=2 take v=u-x